Harry averaged 70 for four straight games, and then disaster struck because after the fifth game his average dropped all the way to 60. What was his score in the fifth game?
Answer by Cromlix(4381)

Total for first 4 games = 4 * 70 = 280
Total for the 5 games = 5 * 60 = 300.
Therefore the score for his 5th game
was 20 (300 - 280)